# Learning Neural Differential Algebraic Equations via Operator Splitting

### v2

@@ V-A2 Model Construction
In the system described by Eq. (12), the differential states are column heights $x_{1}$ and $x_{2}$ and the algebraic states are the volumetric flows $y_{1}$ and $y_{2}$. The input flow is treated as an exogenous control input, i.e., $u=y_{\text{in}}$. We aim to approximate the area-height profile of Tank #1 as well as construct a surrogate for the behavior of the manifold such that (i) fluid is conserved and (ii) the fluid column height constraint is satisfied. There are two neural networks that we seek to train: $NN_{1}:\mathbb{R}^{1}\rightarrow\mathbb{R}^{1}$, an approximation of the sought area-height profile $\phi_{1}$, and $NN_{2}:\mathbb{R}^{4}\rightarrow(0,1)$, the surrogate of the behavior of the manifold. Each is implemented as a feed-forward neural network with one hidden layer of 5 nodes with sigmoid activation functions. The output layer of $NN_{1}$ is linear while the output layer of $NN_{2}$ is sigmoidal, constraining the output to $(0,1)$.
@@ V-A3 Optimization Problem
$\begin{split}\underset{\theta_{1},\theta_{2}}{\text{min}}&~{}~{}\lambda_{1}\sum_{k=1}^{N}||x^{(k)}-\hat{x}^{(k)}||_{2}^{2}+\lambda_{2}\sum_{k=1}^{N}||y^{(k)}-\hat{y}^{(k)}||_{2}^{2}\\ \text{s.t.~{}~{}}&\{x^{(k+1)},y^{(k+1)}\}\\ ~{}&=\operatorname{DAESolve}\left(f,h,\{x^{(k)},y^{(k)},u^{(k)}\};\{\theta_{1},\theta_{2}\}\right),\\ \end{split}$ (13)
@@
$\begin{split}f&:\begin{cases}\frac{y_{1}}{3}\\ \frac{y_{2}}{\operatorname{NN}_{1}(x_{2};\theta_{1})},\\ \end{cases}\\ h&:\begin{cases}u\operatorname{NN}_{2}(x_{1},x_{2},y_{1},y_{2};\theta_{2})\\ u\left(1-\operatorname{NN}_{2}(x_{1},x_{2},y_{1},y_{2};\theta_{2})\right),\\ \end{cases}\\ \end{split}$ (14)
@@ V-A4 Results
The time series data for the tank heights and volumetric flows are given in Figs. 3(a) and 3(b) respectively. The trained model is able to accurately reproduce the time series data seen in training. Additionally, it is able to accurately learn the model of the unknown area-height profile of Tank #2. Figure 4(a) depicts the ground-truth area-height profile and the model’s reconstruction. Furthermore, to demonstrate the generalization, the trained model is tested on an unseen inlet mass flow rate; $u(t)=\frac{1}{2}+\frac{1}{4}\sin{\left(\frac{t}{100}\right)}$. The resulting time series of volumetric flows from the trained model rollout is compared with the ground-truth conditions in Fig. 4(b). Table-I gives the Mean Squared Error (MSE) of the trained neural DAE model.
@@
Extrapolation to a new flow condition demonstrates the ability to (i) endogenize the behaviors of the components in the network and their interaction physics and (ii) disambiguate between data trends and mechanistic relationships contained in the network.
@@
Mean Square Error
@@ V-B1 Problem Description
$\begin{split}\frac{dx_{1}}{dt}&=\frac{1}{\phi_{1}(x_{1})}\left(y_{1}-y_{3}\right)\\ \frac{dx_{2}}{dt}&=\frac{1}{\phi_{2}(x_{2})}\left(y_{2}\right)\\ \frac{dx_{3}}{dt}&=\frac{1}{\phi_{3}(x_{3})}\left(y_{3}-y_{4}\right)\\ \frac{dx_{4}}{dt}&=\frac{1}{\phi_{4}(x_{4})}\left(y_{4}-y_{0}\right)\\ \end{split}$ (15)
$0=y_{0}-p(x_{1},x_{4})\quad$ $0=y_{0}-y_{1}-y_{2}$
$0=y_{3}-\alpha_{1}\sqrt{x_{1}}\quad$ $0=y_{4}-\alpha_{2}\sqrt{x_{3}}$
$0=x_{1}-x_{2}\quad$
@@ V-B2 Optimization Problem
$\begin{split}f&:\begin{cases}\frac{y_{1}-y_{3}}{\phi_{1}(x_{1})}\\ \frac{y_{2}}{\phi_{2}(x_{2})}\\ \frac{y_{3}-y_{4}}{\phi_{3}(x_{3})}\\ \frac{y_{4}-y_{0}}{\phi_{4}(x_{4})},\\ \end{cases}\\ h&:\begin{cases}\operatorname{NN}_{1}(x_{1},x_{4};\theta_{1})\\ y_{0}\operatorname{NN}_{2}(x_{2},x_{3};\theta_{2})\\ y_{0}(1-\operatorname{NN}_{2}(x_{2},x_{3};\theta_{2}))\\ \alpha_{1}\sqrt{x_{1}}\\ \alpha_{2}\sqrt{x_{3}}.\\ \end{cases}\end{split}$ (17)
@@ V-B3 Results
Figure 6(a) details the time history of the four tank heights for both the ground truth data and the open-loop model rollout. At these process conditions, the pump delivers an oscillatory flow to the manifold, which our model is able to capture. Additionally, the tuned model can infer the values for the discharge coefficients: for the case of zero noise, the recovered values are $\alpha_{1}=0.1027$ and $\alpha_{2}=0.1024$. To demonstrate generalizability, the trained model is deployed on an unseen area-height tank profile; for this, we redefine $\phi_{1}(x_{1})=1.0$. In doing so, the system’s limit cycle is destroyed, and the true dynamics approach a steady state. Figure 6(b) demonstrates this extrapolation for the same time horizon. The reconstructed time series data matches the behavior of the unseen data: there exist damped oscillations that decay to a steady state value. Table-II shows the MSE for the neural DAE models trained for the tank network problem. Noise tolerance is an important consideration in the evaluation of system identification tasks, especially in settings with nonlinearities and feedback mechanisms. For this example, we investigate noise tolerance by introducing Gaussian white noise. Figures 7(a) and 7(b) show similar training and extrapolation tasks for the Signal-to-Noise Ratio (SNR) of $20$ dB. Despite the significant noise, the learned dynamics are faithful to the underlying physics, as demonstrated in the extrapolation task, and result in an MSE of $027e-$ 02.
@@ VI CONCLUSIONS AND FUTURE WORKS
This work presents a framework for the data-driven modeling of Differential Algebraic Equations (DAE). Specifically, we introduce a novel Operator Splitting (OS) integration method for Neural Differential Algebraic Equations (NDAE). Inspired by Lie-Trotter operator splitting, we achieve this by splitting the integration of system states into two sequential sub-tasks: first, updating algebraic states via a differentiable surrogate of an algebra solver, followed by a conventional temporal integration step for a system of Ordinary Differential Equations.
The effectiveness of the proposed method is demonstrated by (i) modeling networked dynamical systems with underlying constraints and conservation relationships and (ii) showing generalization to unseen process conditions or component variation. We demonstrate that the proposed OS-based NDAE method is robust to noise and possesses the extrapolation ability to (i) learn the behaviors of the DAE system components and their interaction physics and (ii) disambiguate between data trends and mechanistic relationships contained in the DAE system.
While the methods and example application problems were successful in this work, future research and development are needed in several areas. First, a central assumption of our method is that the underlying dynamics of the systems of interest are sufficiently smooth, i.e., not containing jump discontinuities or shocks, such that the underlying dynamics are non-stiff. As presented, our method leverages an explicit operator splitting method to decouple the evolution of the differential and algebraic states. To increase the ability of the method to handle stiff dynamics or discontinuities, moving to an implicit integration scheme may be warranted. Second, further analysis is needed to investigate the interplay between the modeling task and the index of the underlying DAE (for example, higher-order DAEs that are characteristic of design or parameter optimization problems). Lastly, we have made no claims about computational scalability or partial observability, both of which are critical for modeling real-world complex systems with hundreds of degrees of freedom. We reserve these topics for future work.

### v3

@@ V-A2 Model Construction
In the system described by Eq. (10), the differential states are column heights $x_{1}$ and $x_{2}$, and the algebraic states are the volumetric flows $y_{1}$ and $y_{2}$. The input flow is treated as an exogenous control input, i.e., $u=y_{\text{in}}$. We aim to approximate the area-height profile of Tank #1 as well as construct a surrogate for the behavior of the manifold such that (i) fluid is conserved and (ii) the fluid column height constraint is satisfied. There are two neural networks that we seek to train: $NN_{1}:\mathbb{R}^{1}\rightarrow\mathbb{R}^{1}$, an approximation of the sought area-height profile $\phi_{1}$, and $NN_{2}:\mathbb{R}^{4}\rightarrow(0,1)$, an approximation of the flow manifold connecting two tanks. Each is implemented as a feed-forward neural network with one hidden layer of 5 nodes with sigmoid activation functions. The output layer of $NN_{1}$ is linear while the output layer of $NN_{2}$ is sigmoidal, constraining the output to $(0,1)$.
@@ V-A3 Optimization Problem
$\begin{split}\underset{\theta_{1},\theta_{2}}{\text{min}}&~{}~{}\lambda_{1}\sum_{k=1}^{N}||x^{(k)}-\hat{x}^{(k)}||_{2}^{2}+\lambda_{2}\sum_{k=1}^{N}||y^{(k)}-\hat{y}^{(k)}||_{2}^{2}\\ \text{s.t.~{}~{}}&\{x^{(k+1)},y^{(k+1)}\}\\ ~{}&=\operatorname{DAESolve}\left(f,h,\{x^{(k)},y^{(k)},u^{(k)}\};\{\theta_{1},\theta_{2}\}\right).\\ \end{split}$ (11)
@@
$\begin{split}f:\begin{cases}\frac{y_{1}}{3}\\ \frac{y_{2}}{\operatorname{NN}_{1}(x_{2};\theta_{1})}\\ \end{cases},\quad h:\begin{cases}u\operatorname{NN}_{2}(x_{1},x_{2},y_{1},y_{2};\theta_{2})\\ u\left(1-\operatorname{NN}_{2}(x_{1},x_{2},y_{1},y_{2};\theta_{2})\right)\\ \end{cases}.\\ \end{split}$ (12)
@@ V-A4 Results
The time series data for the tank heights and volumetric flows are given in Figs. 3(a) and 3(b) respectively. The trained model accurately reproduces the time series data. Additionally, it is able to accurately learn the model of the unknown area-height profile of Tank #2. Figure 4(a) depicts the ground-truth area-height profile and the model’s reconstruction. Furthermore, to demonstrate the generalization, the trained model is tested on an unseen inlet mass flow rate; $u(t)=\frac{1}{2}+\frac{1}{4}\sin{\left(\frac{t}{100}\right)}$. The resulting time series of volumetric flows from the model rollout is compared with the ground-truth conditions in Fig. 4(b). Table-I gives the Mean Squared Error (MSE) of the trained neural DAE model.
@@
Extrapolation to a new flow condition demonstrates the ability to (i) endogenize the behaviors of the components in the network and their interaction physics and (ii) disambiguate between data trends (such as correlations arising from specific operating regimes) and mechanistic relationships contained in the network.
@@
Mean Square Error (MSE)
@@ V-B1 Problem Description
$\begin{split}\frac{dx_{1}}{dt}&=\frac{1}{\phi_{1}(x_{1})}\left(y_{1}-y_{3}\right),\quad\frac{dx_{2}}{dt}=\frac{1}{\phi_{2}(x_{2})}\left(y_{2}\right),\\ \frac{dx_{3}}{dt}&=\frac{1}{\phi_{3}(x_{3})}\left(y_{3}-y_{4}\right),\quad\frac{dx_{4}}{dt}=\frac{1}{\phi_{4}(x_{4})}\left(y_{4}-y_{0}\right),\\ \end{split}$ (13)
00 0 $=y_{0}-p(x_{1},x_{4}),\quad 0=y_{0}-y_{1}-y_{2},$
00 0 $=y_{3}-\alpha_{1}\sqrt{x_{1}},\quad 0=y_{4}-\alpha_{2}\sqrt{x_{3}},\ \quad 0=x_{1}-x_{2}.$
@@ V-B2 Optimization Problem
$\begin{split}f:\begin{cases}\frac{y_{1}-y_{3}}{\phi_{1}(x_{1})}\\ \frac{y_{2}}{\phi_{2}(x_{2})}\\ \frac{y_{3}-y_{4}}{\phi_{3}(x_{3})}\\ \frac{y_{4}-y_{0}}{\phi_{4}(x_{4})}\\ \end{cases},\quad h:\begin{cases}\operatorname{NN}_{1}(x_{1},x_{4};\theta_{1})\\ y_{0}\operatorname{NN}_{2}(x_{2},x_{3};\theta_{2})\\ y_{0}(1-\operatorname{NN}_{2}(x_{2},x_{3};\theta_{2}))\\ \alpha_{1}\sqrt{x_{1}}\\ \alpha_{2}\sqrt{x_{3}}\\ \end{cases}\end{split}$ (15)
@@ V-B3 Results
Figure 6(a) shows open-loop trajectories of the four tank heights. At these process conditions, the pump delivers an oscillatory flow to the manifold, which our model is able to capture. Additionally, the tuned model can infer the values for the discharge coefficients: for the case of zero noise, the recovered values are $\alpha_{1}=0.1027$ and $\alpha_{2}=0.1024$. To evaluate generalization, we test the trained model on an unseen area–height relationship that was not used during training. Specifically, while the model is trained on one tank geometry with a particular cross-sectional area profile (training case), we deploy it on a different tank with a modified area–height profile (validation case); for this, we redefine $\phi_{1}(x_{1})=1.0$. In doing so, the system’s limit cycle is destroyed, and the true dynamics approach a steady state. Figure 6(b) demonstrates this extrapolation for the same time horizon. The reconstructed time series data matches the behavior of the unseen data: there exist damped oscillations that decay to a steady state value. Table-II shows the MSE for the neural DAE models trained for the tank network problem.
Noise tolerance is an important consideration in the evaluation of system identification tasks, especially in settings with nonlinearities and feedback mechanisms. For this example, we investigate noise tolerance by introducing Gaussian white noise. Figures 7(a) and 7(b) show similar training and extrapolation tasks for the Signal-to-Noise Ratio (SNR) of $20$ dB. Despite the significant noise, the learned dynamics are faithful to the underlying physics, as demonstrated in the extrapolation task, and result in an MSE of $027e-$ 02.
@@ VI CONCLUSIONS AND FUTURE WORKS
This work presents a data-driven framework for modeling differential algebraic equations (DAEs) using a novel operator splitting (OS) method for Neural DAEs. Inspired by Lie-Trotter splitting, our approach sequentially updates algebraic states via a differentiable surrogate and integrates differential states using an ODE solver. The effectiveness of the proposed method is demonstrated by (i) modeling networked dynamical systems with algebraic constraints and (ii) showing generalization to unseen process conditions or component variation. Furthermore, demonstrate robustness to noise and extrapolation to (i) learn the behaviors of the DAE system components and their interaction physics and (ii) disambiguate between data trends and mechanistic relationships contained in the DAE system.
Future research is needed in several areas. First, a baseline comparison with standard neural ODEs or PINNs would help quantify the gains of our approach. Second, a key assumption of the current formulation is that the dynamics are sufficiently smooth and non-stiff. To extend applicability to stiff or discontinuous dynamics, implicit integration schemes and adaptive solvers should be explored. Third, extending the framework to higher-index DAEs (characteristic of many optimal design and control applications) warrants further theoretical and numerical investigation. Lastly, this work has not addressed computational scalability or partial observability, both of which are critical for modeling large-scale systems with hundreds of degrees of freedom. We reserve these directions for future work.